OpenStudy (yisselluna):

Solve the following system of equations. 3x + 2y - 5 = 0 x = y + 10 ANSWER: {( a0)}

OpenStudy (andrewishere):

Would you like my help on this one too?

OpenStudy (yisselluna):

yes plsss

OpenStudy (andrewishere):

Ok so this one is a little different. This time we have to solve the system of equations. Since the second equation is already in terms of x, let's substitute that into the top equation. Now, I am sure you are a pro in substituting now as of our last problem ;) So I will give you the quick and simple version: 3(y+10)+2y-5=0 Now distribute and combine like terms: 3y+30+2y-5 = 0 5y+25 = 0 Now put 5y+25 = 0 into terms of y. 5y = -25 y = -5 (We need this)

OpenStudy (andrewishere):

Now that we know that y = -5 (keep this in mind, we will need this to complete the problem), substitute -5 into the second equation, x=y+10 x = y+10 x = (-5) + 10 x = 5 Now we know both the x and the y values. x = 5 y = -5 Put this into coordinate form (x,y) So the answer is (5,-5)

OpenStudy (andrewishere):

Hope this helped

OpenStudy (yisselluna):

thank u so muchh!!!

OpenStudy (andrewishere):

No problem